Detailed proof please.. 1. Determine whether the following statements are true or false. If one is true, provide a proof. If one is false, provide
Answer #1
  • FALSE. Let f(x) = \frac1x . Then \int_1^\infty (f(x))^2\,\text dx =\int_1^\infty \frac1{x^2}\,\text dx = \left[ -\frac 1x \right ]_1^\infty =1 converges but \int_1^\infty f(x)\,\text dx =\int_1^\infty \frac1x\,\text dx = \left[ \ln x \right ]_1^\infty =\infty does not.
  • TRUE. Since lim f(0) = 0 T- there exists a R>1 such that f(x)<1 for x\geqslant R . For x\geqslant R we have (f(x))^2\leqslant f(x) . Then,
    \int_1^\infty (f(x))^2 \,\text dx = \int_1^R (f(x))^2 \,\text dx+ \int_R^\infty \underbrace {(f(x))^2}_{\leqslant f(x)} \,\text dx
    \leqslant \int_1^R (f(x))^2 \,\text dx+ \int_R^\infty f(x) \,\text dx
    <\infty .
  • FALSE. Let f such that f(n) = 1 for n \in \mathbb N . Let f\left(n \pm \frac{1}{n^2}\right) =0 and extend f linearly between the intervals \left[ n- \frac{1}{n^2},n \right ] and \left[ n,n + \frac{1}{n^2}\right ] . Put f(x) = 0 elsewhere. The function looks like triangles of height 1 with base \frac{2}{n^2} centered at n . Then
    \int_1^\infty f(x) \,\text dx = \frac 14 + \sum_{n=2}^\infty \frac {\text{base}\times\text{height}}{2} = \frac 14 + \sum_{n=2}^\infty \frac{1}{n^2}<\infty converges
    but
    \sum_{n=1}^\infty f(n) = \sum_{n=1}^\infty 1 = \infty does not.
  • FALSE. Let f(x) = |\sin(\pi x)| . Then \int_1^\infty f(x) \,\text dx = \int_1^\infty |\sin (\pi x)| \,\text dx = \infty but \sum_{n=1}^\infty f(n) = \sum_{n=1}^\infty |\sin (\pi n)| = 0 .